The electric field has a constant value of 4.1 x 104 V/m and is directed downward. The field is the same everywhere. The potential at a point P within this region is 169 V. Find the potential at the following points: (a) 9.4 x 10-3 m directly above P, (b) 3.8 x 10-3 m directly below P, (c) 6.2 x 10-3 m directly to the right of P.
Answer #1

Electric field E = 41000 V / m

Potential at point P is V = 169 V

(a).  Potential at 9.4*10^-3 m directly above P is V ' = V + ( Ed )

Where d = 9.4*10^-3 m

Substitue values we get V ' = 169 + 385.4

                                           = 554.4 volt

b)Potential at 3.8*10^-3 m directly below P is V " = V - Ed'

Where d ' = 3.8*10^-3 m

Substitue values we get V " = 169 -155.8

                                           = 13.2 volt

c) potential at 6.2*10^-3 m directly to the right of P is V"' = V

= 169 volt

I hope help you !!